OpenStudy (akohl):

@Abhisar A spring is compressed to a displacement of -3.9 m from equilibrium. Then the spring is released and allowed to stretch to a displacement of -1.5 m from equilibrium. The spring constant is 11 N/m. What best describes the work involved as the spring stretches? 71 J of work is performed on the spring. 140 J of work is performed on the spring. The spring performs 71 J of work. The spring performs 140 J of work

OpenStudy (abhisar):

What do u think it should be ?

OpenStudy (akohl):

I got 71 but that makes sense for both options with 71

OpenStudy (abhisar):

u got 71 for which part and how ?

OpenStudy (akohl):

I did .5*11*3.9^2 -.5*11*1.5^2

OpenStudy (abhisar):

When spring is compressed then some work is done on it. When spring is stretched then spring does some work.

OpenStudy (abhisar):

When we compress a spring, then the amount of work done on it gets stored as potential energy.

OpenStudy (abhisar):

PE=0.5*k*x^2

OpenStudy (abhisar):

In this case it will be equal to 0.5*3.9^2*11=83.65

OpenStudy (abhisar):

Now when the spring is released, uses the stored potential energy in it and does some work. In this case the spring is allowed to move upto 1.5 from equillibrium.

OpenStudy (abhisar):



OpenStudy (abhisar):

So the work done by the spring will be potential energy used which in turn will be final potential - initial potential energy.

OpenStudy (abhisar):

0.5*11*1.5^2 - 0.5*11*3.9^2 = -71

OpenStudy (abhisar):

negative sign shows that the work is done by the spring.

OpenStudy (akohl):

Awesome that makes sense now thanks

OpenStudy (abhisar):

you're most welcome !